A restaurant manager bought 20 packages
of bagels. Some packages contained 6 bagels
each, and the rest contained 12 bagels each.
There were 168 bagels in all. How many
packages of 12 bagels did the manager buy?

Answer by mananth(16949)

6 bagel packages ------- x numbers
12 bagel packages ----------y numbers
x+y =20
6x+12y=168
x + y = 20 .............1
6x + 12y = 168 .............2
Eliminate y
multiply (1)by -12
Multiply (2) by 1
-12x -12y = -240
6x 12y = 168
Add the two equations
-6 x = -72
/ -6
x = 12.00 OR 12
plug value of x in (1)
1 x + 1 y = 20
12 + 1 y = 20
1 y = 20-12
1 y = 8

6 pack 12 numbers
12 pack 8 umbers
m.ananth@hotmail.ca